A bicycle path is 7 miles long. A man walks from one end at the rate of 3 mph. At the same time, a friend bicycles from the other end, traveling at 25 mph. In how many minutes will they meet?
Answer by stanbon(75887)

A bicycle path is 7 miles long. A man walks from one end at the rate of 3 mph. At the same time, a friend bicycles from the other end, traveling at 25 mph. In how many minutes will they meet?
----------------
Walker DATA:
rate = 3 mph ; distance = x miles ; time = x/3 hrs
-------------------------
Bicyclist DATA:
rate = 25 mph ; distance = 7-x miles ; time = (7-x)/25 hrs.
--------------------
Equation:
time = time
x/3 = (7-x)/25
25x = 21-3x
28x = 21
x = 21/28 miles
-----
Time for the walker = x/3 hrs = (21/28)/3 hrs = 7/28 = 1/4 hr. = 15 minutes
They will meet in 15 minutes.
====================
cheers,
Stan H.